Titanium crystallizes in a body-centered cubic unit cell with an edge length of 3.306 Å. The density of titanium is 4.401 g/cm3. Use these data to calculate Avogadro’s number.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Afridi
 
Volume of unit cell =  (3.306)³
 = 36.13 A°³ = 36.13 * 10^-24 A°
 = 3.613 * 10^-23 A°
 
Mass of unit cell = 3.613 * 10^-23 * 4.401 = 1.590 * 10^-22 gm
 
BCC unit cell consisits of 2 (Ti) atoms
Hence mass of one atom = 1.590 *10^-22/2
= 7.590 * 10^-23
 
Atomic weight of Ti(47.88)
 
Hence Avogadro number N = 47.88/ 7.95 * 10^-23
N = 6.023 * 10^23
 
Regards
Arun (askIITians forum expert)